1)an ideal gas is taken through a process in which the pressure and volume are related as p=kv.show that molar heat capacity of the gas for the process is given by c = Cv + R/2

(Cv - molar heat capacity at constant volume)

2)an ideal gas with Cp/Cv = is taken through a process in which pressure and volume vary as p = av^b..find the value of b for which the specific heat capacity for the process is zero..

please post with complete solutions


Grade:upto college level

1 Answers

Samuel Garry
5 years ago
Sol. P = KV ⇒ nRT/V = KV ⇒ RT = KV^2 ⇒ R ∆T = 2KV ∆U ⇒ R∆T/2KV = dv dQ = du + dw ⇒ mcdT = C base VdT + pdv ⇒ msdT = C base VdT + PRdF/2KV ⇒ ms = C base V + RKV/2KV ⇒ C base p + R/2